Can someone help me solve this?
Type of Problem: Equations, Inequalities, and Absolute Value
Solve the inequality and graph the solution.
|2d-2|<10.8

Answer by waynest(281)

|2d - 2| < 10.8
2d - 2 = 10.8
2d - 2 + 2 = 10.8 + 2
2d = 12.8
2d/2 = 12.8/2
d = 6.4
d < 6.4